the bumper on a 1200kg car is tested by driving it into a brick wall. The bumper is equivalent to a spring with spring constant 6 x 10^6 N/m and it compresses 3.40 cm to stop the car. if all the energy is absorbed by the bumper elastically, what was the car’s speed before it hit

3. The bumper on a 1200kg car is tested by driving it into a brick wall. The bumper is ivalent to a spring with spring constant 6 x 10° N/m and it compresses 3.40cm to stop the car. If all the energy is absorbed by the bumper elastically, what was the cars speed before it hit?
Answer #1

Energy stored in the bumper = 0.5*6*10^6*0.034^2= 3468

so, 3468 = 0.5*1200*v^2

v= 2.4m/s